Circular No. 8223

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2003 T3 Vello Tabur, Australian Capital Territory, reports his discovery of a

somewhat-condensed comet with a 30′′ coma on unfiltered CCD images taken with a 140-mm f/2.8 camera lens on Oct. 14.481 UT. T. Lovejoy (Thornlands, Queensland, 0.16-m reflector) reports that a CCD exposure taken on Oct. 15.5 shows a 0′.7 coma of total mag 11.6 and a fan-shaped tail ∼ 1′ long in p.a. 90^{o}. Available astrometry:

2003 UT  α_{2000}  δ_{2000}  Mag. Observer
Oct. 14.5750 20^{h}00^{m}40.^{s}09 −57^{o}23′07′′.4 11.7 Tabur
14.5888 20 00 40.18 −57 22 51.1  _{”}
14.5975 20 00 40.35 −57 22 42.0  _{”}
14.6046 20 00 40.59 −57 22 33.7  _{”}
14.6123 20 00 40.78 −57 22 23.5  _{”}
14.6235 20 00 40.96 −57 22 12.1  _{”}
15.37314 20 00 55.12 −57 07 56.4  Broughton
15.37791 20 00 55.22 −57 07 50.9  _{”}
15.39804 20 00 55.57 −57 07 27.5 12.0  _{”}
15.46418 20 00 56.73 −57 06 11.7 14.1 Higgins
15.47361 20 00 56.92 −57 06 00.6 14.3  _{”}
15.48303 20 00 57.10 −57 05 49.7 14.2  _{”}

V. Tabur (Wanniassa, A.C.T.). 0.28-m reflector + CCD. J. Broughton (Reedy Creek, Qld.). 0.25-m reflector + CCD. 1′ coma with

a 3′ dust tail in p.a. 200^{o}. D. Higgins (Ngunnawal, A.C.T.). 0.25-m reflector + CCD.

1937 UB (HERMES) B. A. Skiff, Lowell Observatory, reported a bright (R = 14.4) near-

earth-object candidate from LONEOS images taken on Oct. 15.3 UT. After posting on the NEO Confirmation Page, follow-up observations were made by J. Young at Table Mountain. T. Spahr, Minor Planet Center, then recognized NEAT observations from Oct. 5 and identified the object as the long-lost 1937 UB (Hermes). The available astrometry (including prerediscovery images back to Aug. 26), orbital elements, and an ephemeris appear on MPEC 2003-T74. The object has made 31 revolutions about the sun since its discovery.

2003 October 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green